A merchant has coffee worth $20 a pound that she wishes to mix with 50 pounds of coffee worth $50 a pound to get a mixture that can be sold for $40 dollars a pound. How many pounds of the $20 coffee should be used?
Answer by htmentor(1343)

Let x = the number of lbs of $20 coffee
Since there are 50 lbs of $50 coffee, the final mixture will contain x+50 lbs
We want the mixture to sell for $40 per pound, so we can write
40(x+50) = 50*50 + 20x
Solve for x:
40x + 2000 = 2500 + 20x
20x = 500
x = 25
Therefore, she should use 25 lbs of $20 coffee